Genetics Test- please help! :)  Am I correct?

1: In a human population, the genotype frequencies at one locus are 0.5 AA, 0 The frequency of the A allele is: 4 Aa, and 0.1 aa fA (o,4) t
Answer #1

Yes absolutely!!

f (A) = f (AA) +1/2 (Aa)

0.5 + 1/2 (0.4)

0.5 + 0.2 = 0.7

or you can use this formula =A2 + 2Aa + a2